Plastic parts produced by an injection-molding operation are checked for conformance to specifications. Each tool contains 13 cavities in which parts are produced, and these parts fall into a conveyor when the press opens. An inspector chooses 2 part(s) from among the 13 at random. Three cavities are affected by a temperature malfunction that results in parts that do not conform to specifications Round your answers to four decimal places. (a) What is the probabilty that the inspector finds exactly one nonconforming part? (b) What is the probability that the inspector finds at least one nonconforming part?
Answer #1

Here, X:Number of non conforming parts obtained

as P(Getting a non conforming part) =3/13=0.2308(approx)

a)P(X=1)=2C1*0.2308*(1-0.2308)=0.3551(a Ans)

b) P(X>=1)=P(X=1)+P(X=2)=0.3551+2C2*0.2308^2*(1-0.2308)^0=0.3551+0.0533=0.4084(b Ans)